OpenStudy (anonymous):

PLZ HELP ME Solve 144⋅x2 - 36 = 0 by using square roots

OpenStudy (anonymous):

@kirbykirby can u help me again plz ur awesome??

OpenStudy (anonymous):

is this right? http://www.tiger-algebra.com/drill/144$x~2-36=0/

OpenStudy (anonymous):

@tkhunny @uri can u help me plzzz

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (tkhunny):

I don't recommend "using square roots", but have you considered adding 36? That would be the place to start.

OpenStudy (anonymous):

its what the problem says out of the book sir or ma'am

OpenStudy (tkhunny):

I get it. Thus my question. Have you considered adding 36? That would be the place to start.

OpenStudy (kirbykirby):

\[ 144x^2-36=0\\ 144x^2=36\\ x^2=\frac{36}{144}\\ x^2=\frac{1}{4}\\ x=\pm\sqrt{\frac{1}{4}}\\ x=\pm\frac{1}{2}\]

OpenStudy (kirbykirby):

The method on the site is correct too, although a bit longer.